OpenStudy (anonymous):

Prove or Disprove: Given k; n belong to N If k/n! then k/i for some 0 <= i <= n.

OpenStudy (anonymous):

consider \(n=0\). clearly \(n!=1\); if \(k|1\) it follows that \(k=1\) but \(0\le 1\not\le 0\)

OpenStudy (anonymous):

but 0 does not belong to N

OpenStudy (anonymous):

so that disproves it in the case that \(0\in\mathbb{N}\)

OpenStudy (anonymous):

in our class 0 does not belong to N

OpenStudy (anonymous):

so do you think that the proof will be by induction?

OpenStudy (anonymous):

take \(n=4\) so \(n!=24\); clearly \(12|n!\) so we can take \(k=12\). it's easy to check, however, that \(12\) does not divide any of \(0,1,2,3,4\)

OpenStudy (zarkon):

go smaller... 6|3!

OpenStudy (anonymous):

\(n=4\) is just a sufficiently easy yet dramatic counterexample that I found it more useful but you're correct, \(n=3\) with \(k=6\) also works

OpenStudy (zarkon):

so i can be 0

OpenStudy (zarkon):

\[0\le i\le n\]

OpenStudy (zarkon):

for any m in the natural numbers m|0 since 0=0m

OpenStudy (anonymous):

i got it, i was dividing!so i tought is true. Thanks!

OpenStudy (zarkon):

if \(i\) can be zero then the problem is booring

OpenStudy (zarkon):

here \(m\in\mathbb{N}=\{1,2,3,\ldots\}\) so no problem

OpenStudy (zarkon):

since \(m\in\mathbb{N}\) then \(m|0\) since \(m0=0\)

OpenStudy (anonymous):

oops, misread

OpenStudy (anonymous):

then @Zarkon is correct, so it follows the \(i=0\) possibility makes this trivially true for all \(k,n\)

OpenStudy (zarkon):

you are not looking at it correctly. for this problem \(k\in\mathbb{N}\) so it can't be zero then I can find an integer (in this case 0) such that k0=0

OpenStudy (zarkon):

like i said...this problem is boring if \(i\) can take the value 0

OpenStudy (anonymous):

oh so it is true since i equal to 0 but how can we proof it then?